OpenStudy (anonymous):

What is the range of the graph of y = (x - 7)^2 + 2? y greater than or equal to 7 y less than or equal to 7 y less than or equal to 2 y greater than or equal to 2

OpenStudy (aum):

y = (x - 7)^2 + 2 \((x-7)^2 \ge 0\) (because squaring turns negative to positive). Therefore, \(y \ge 2\)

jimthompson5910 (jim_thompson5910):

If we have the equation \[\Large y = a(x - h)^2 + k\] where a > 0, then the range is \(\Large y \ge k\) ------------------------------------------------------- If we have the equation \[\Large y = a(x - h)^2 + k\] where a < 0, then the range is \(\Large y \le k\)

OpenStudy (aum):

y = (x - 7)^2 + 2 This is vertical, upright parabola with vertex at (7, 2). In a vertical upright parabola, the vertex is the lowest point. Therefore, y will always be greater than or equal to the y-coordinate of the vertex. \(y \ge 2\)